OpenStudy (anonymous):

4y+19=x, 3y-x=-13 solve using substitution

OpenStudy (anonymous):

4y+19=x which mean : x = 4y + 19 subtitute into 2nd equation 3y-x=-13 3y-(4y+19)=-13 3y-4y-19=-13 which y to the RHS -19+13=y therefore y=-6 then substitute this value into the equation and get your x cheers :)

OpenStudy (anonymous):

Thank u I got two more

OpenStudy (anonymous):

I need help with x now lol

OpenStudy (anonymous):

oh, try to substitute the value of y just now into the 1st equation :D

OpenStudy (anonymous):

show ur working here, i'll help u :)

OpenStudy (anonymous):

I did that I got 4(-6)+19=x

OpenStudy (anonymous):

then, just proceed :D

OpenStudy (anonymous):

multiply 4 with -6

OpenStudy (anonymous):

-18

OpenStudy (anonymous):

nope, try again :p

OpenStudy (anonymous):

I mean -24

OpenStudy (anonymous):

lol, ok, now add with 19 :)

OpenStudy (anonymous):

-43

OpenStudy (anonymous):

add ! ><

OpenStudy (anonymous):

seems like u're confusing negative part n positive part, ill try to draw the concept so u can understand better :D

OpenStudy (anonymous):

Is it -5

OpenStudy (anonymous):



OpenStudy (anonymous):

remmber that any valu u add, the direction will go towards +ve side, and minus will go to the -ve side, so, when u got -ve value, and be added with something, that value will get closer to 0, and if -ve value is minus with something, it will get further from 0

OpenStudy (anonymous):

yeah, u're doing right for -5 :D

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

Now I need elimination

OpenStudy (anonymous):

For 3x+2=17 4x-2y=4

OpenStudy (anonymous):

u better to open a new question cause i got class later :)

OpenStudy (anonymous):

Ok